If 237 g of H2 is mixed with N2 at a pressure of 152,000 torr and a temperature of 400°C, what volume of N2 gas is needed for the reaction?
1
Expert's answer
2020-04-13T02:34:03-0400

237 g of H2 is mixed with N2 at a pressure of 152,000 torr and a temperature of 400°C, what volume of N2 gas is needed for the reaction?

N2 + 3 H2 = 2 NH3

The amount of H2 is n(H2) = 237 g / 2 g/mole = 118.5 moles.

Therefore, the amount of N2 needed is n(N2) = 118.5 / 3 = 39.5 moles.

From the gas law p*V=n*R*T,

the volume of N2 needed is V(N2) = n(N2)*R*T/p.

T = 400°C +273 = 673 K.

p = 152,000 torr / 760 = 200 atm = 200*101325 Pa = 20265000 Pa.

The volume of N2 is V(N2) = 39.5*8.31*673/20265000 = 0.0109 m3.

The volume of N2 needed is 10.9 L.